OpenStudy (anonymous):

solve using the quadratic formula. I'm trying to graduate, so guys or girls please help me. 3x^2-8x+2=0

OpenStudy (john_es):

The solution for the general equation, \[ax^2+bx+c=0\] is \[x=\frac{b\pm\sqrt{b^2-4ac}}{2a}\] You can do it from this point.

OpenStudy (anonymous):

I'm not understanding this formula.

OpenStudy (john_es):

Compare the general equation with your equation. You will see that, \[a=3\\ b=8 \\ c=2\] Then substitute in the solution, \[x=\frac{-8\pm\sqrt{8^2-4\cdot3\cdot 2}}{2\cdot 3}\] You will have two solutions, one for each sign of the square root, \[x=\frac{-8-\sqrt{8^2-4\cdot3\cdot 2}}{2\cdot 3}\\\]

OpenStudy (john_es):

And \[x=\frac{-8+\sqrt{8^2-4\cdot3\cdot 2}}{2\cdot 3}\]

OpenStudy (anonymous):

?

OpenStudy (anonymous):

In the formula it should be [-b. +- sqrt(b^2-4ac)]/2a first - missing Example worked properly

OpenStudy (anonymous):

right, but how do i work these types of problems?

OpenStudy (anonymous):

re-write them so that they fit the format ax^2 + bx + c = 0 then use formula for solving quadratic equations, as given above note that there will be two solutions because of the +- sign in the formula. For a real problem, one of these solutions will make sense, the other will not. If the sqrt has a negative b^2 - 4 a c, then you will have tow "imaginary" solutions, with format x = something +- i somethingelse

OpenStudy (anonymous):

I'm sorry but I'm very very lost right now.

OpenStudy (anonymous):

What is the question you have about what has been presented?

OpenStudy (anonymous):

all it says is to put the Approximate answers to the nearest tenth.

OpenStudy (anonymous):

Use the formula, get an answer, round it up or down, depending on whether the decimal part of the answer is greater or less than 0.A5: <0.A5 becomes 0.A rounded down >= 0.A5 becomes 0.(A+1) rounded up e.g., 2.15 = 2.2 2.14 = 2.1

OpenStudy (anonymous):

?

OpenStudy (anonymous):

x = 2.39, 0.28

OpenStudy (anonymous):

oh........ok